# Reinforced Prompt Personalization for Recommendation with Large Language Models

### v1

@@ Abstract.
Designing effective prompts can empower LLMs to understand user preferences and provide recommendations by leveraging LLMs’ intent comprehension and knowledge utilization capabilities. However, existing research predominantly concentrates on task-wise prompting, developing fixed prompt templates composed of four patterns (i.e., role-playing, history records, reasoning guidance, and output format) and applying them to all users for a given task. Although convenient, task-wise prompting overlooks individual user differences, leading to potential mismatches in capturing user preferences. To address it, we introduce the concept of instance-wise prompting to personalize discrete prompts for individual users and propose Reinforced Prompt Personalization (RPP) to optimize the four patterns in prompts using multi-agent reinforcement learning (MARL). To boost efficiency, RPP formulates prompt personalization as selecting optimal sentences holistically across the four patterns, rather than optimizing word-by-word. To ensure the quality of prompts, RPP meticulously crafts diverse expressions for each of the four patterns, considering multiple analytical perspectives for specific recommendation tasks. In addition to RPP, our proposal of RPP+ aims to enhance the scalability of action space by dynamically refining actions with LLMs throughout the iterative process. We evaluate the effectiveness of RPP/RPP+ in ranking tasks over various datasets. Experimental results demonstrate the superiority of RPP/RPP+ over traditional recommender models, few-shot methods, and other prompt-based methods, underscoring the significance of instance-wise prompting for LLMs in recommendation tasks and validating the effectiveness of RPP/RPP+. Our code is available at https://github.com/maowenyu-11/RPP.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Transactions on Information Systems; July 03–05, 2024; Woodstock, NYccs: Computing methodologies Discrete space searchccs: Computing methodologies Multi-agent reinforcement learningccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RSs) are widely applied in real life to offer personalized item recommendations that align with user preference. Recently, large language models (LLMs), such as ChatGPT (Ouyang et al., 2022), GPT4 (Bubeck et al., 2023), PaLM-E (Driess et al., 2023), and LLaMA (Touvron et al., 2023), have demonstrated exceptional capabilities in semantic understanding, intent reasoning, and knowledge utilization (Zhao et al., 2023). These advancements empower LLMs to address recommendation tasks by modeling users’ behavior into natural language, attracting considerable attention (Li et al., 2023c; Geng et al., 2022; Zhang et al., 2023b; Liao et al., 2024).
To unleash the recommendation capabilities of LLMs, a promising solution lies in tailoring prompts for the specific recommendation task (Zhang et al., 2023b; Gao et al., 2023; Geng et al., 2022), beyond fine-tuning the massive parameters (Bao et al., 2023) that can be resource-hungry. Existing methods (Wu et al., 2023; Wang and Lim, 2023; Gao et al., 2023; Hou et al., 2024) have investigated the task-wise prompting through manual engineering or enumeration followed by selection (Deng et al., 2022), dedicated to establishing fixed prompt templates shared across all users for a given recommendation task. To summarize, most of these task-wise prompts for recommendation tasks include the following patterns:
@@
  • History records provides the history records of users’ interaction, which the LLMs can utilize as in-context information for subsequent recommendations (Hou et al., 2024; Wang and Lim, 2023; Liao et al., 2024).
  • Reasoning guidance guides LLMs with essential steps in the reasoning process to make recommendations, incorporating strategies such as chain-of-thought (CoT) (Wei et al., 2022; Liao et al., 2024; Hou et al., 2024).
@@
However, we argue that applying fixed task-wise prompts for all users in a task (Wu et al., 2022; Jin et al., 2023) may fail to fully explore the capability of LLMs to provide personalized recommendations. First of all, the one-fitting-all prompting falls short of accurately capturing users’ varied intentions and preferences, which may result in mediocre recommendation performance. Take Figure 1 as an example, the first user’s preference for science fiction movies is reflected in their long-term watching history with films like Superman and Cocoon. In contrast, the second user’s preference for comedy movies is evident in their short-term watching history with films like Grumpy Old Men and Sister Act. Thus, task-wise prompts, such as instructing LLMs to recommend based on users’ long-term or short-term watching history, is not applicable to all users. Secondly, the performance of LLMs is sensitive to the expression of prompts (Lu et al., 2022; Zhu et al., 2023), fixed prompt templates limit the exploration of the advantages that varying expressions of prompts can bring to the output of LLMs. Therefore, we propose personalizing instance-wise prompts for different users in the recommendation tasks, as shown in Figure 1.
While conceptually appealing, personalizing the discrete prompts for individual users is challenging on prompt optimization. Existing discrete optimization methods (e.g., manual creation (Brown et al., 2020; Reynolds and McDonell, 2021) or heuristical crafting (Jiang et al., 2020; Pan et al., 2023)) are labor-intensive and resource-expensive to tailor prompts for each user. To close this gap, we draw inspiration from reinforcement learning (Kaelbling et al., 1996) to personalize prompts automatically. The basic idea is framing the space of natural language as the action space and formulating the prompt generation as selecting actions based on the user’s personalized information (Deng et al., 2022; Li et al., 2023b), such as the encoded user representation, user interaction history, and user preferences. Such personal information guides the learning of the policy network, personalizing prompts for individual users iteratively. Recent works (Deng et al., 2022; Zhang et al., 2023a) have delved into optimizing prompts using RL. However, these efforts have either involved searching for tokens within an extensive vocabulary or conducting basic editing operations on the prompts, constrained by computational resource limitations and the reliability of the generated prompts. To apply RL for personalizing instance-wise prompts in the recommendation tasks, two fundamental problems are crucial:
Figure 1. Comparison between task-wise prompting and instance-wise prompting for recommendation.
  • How to constrain the search space to enhance efficiency?
@@
To tackle these problems, we introduce Reinforced Prompt Personalization (RPP) to personalize prompts for each user with multi-agent RL (MARL) algorithm (Zhang et al., 2019). We limit the search space by exploring the various expressions of the four patterns at the sentence level, i.e., {role-playing, history records, reasoning guidance, output format}, instead of searching word-by-word in the extensive vocabulary. To ensure prompt quality, we meticulously design the diverse expressions of each pattern based on the nature of recommendation tasks, which can guide LLMs to understand users’ intent and make recommendations from multiple perspectives. The personalized prompts interact with LLMs which function as the environment, to maximize LLMs’ recommending performance which serves as the reward. Specifically, RPP employs multi-agent RL based on the Actor-Critic framework (Konda and Tsitsiklis, 1999), adhering to the Centralized Training with Decentralized Execution (CTDE) paradigm  (Lowe et al., 2017). Each agent is trained with an individual policy network while having access to a shared global state, personalizing a specific prompt pattern. This global state contains information about the personalized prompts and the ranking results generated by LLMs in the current round. The action space is carefully designed with each action defined through collaborative deliberation between human experts and LLMs. Moreover, to improve the flexibility and quality of the action space, we dynamically refine the selected actions with LLMs during the iterative process to enhance scalability, referred to as RPP+.
@@
We assess the effectiveness of our approach in ranking tasks following the setting of LLMRank (Hou et al., 2024). Experimental results demonstrate that RPP/RPP+ can significantly improve the recommendation performance of LLMs, outperforming several traditional recommender models (e.g., BPRMF (Rendle et al., 2009), SASRec (Kang and McAuley, 2018)), few-shot methods (e.g., VQ-Rec (Hou et al., 2023), UniSRec (Hou et al., 2022)), and other prompt-based methods (e.g., manual prompts (Hou et al., 2024), enumeration) across three widely used public datasets(i.e., MovieLens-1M (Harper and Konstan, 2016), Games (Ni et al., 2019), and Lastfm (Cantador et al., 2011)). This highlights the significance of personalizing instance-wise prompts for LLMs to recommend and the superiority of RPP/RPP+ in enhancing the recommendation ability of LLMs. Our key contributions are as follows:
@@
  • We propose personalizing instance-wise prompts for each user in recommendation tasks and introduce a new framework, Reinforced Prompt Personalization (RPP/RPP+), to address it.
  • We employ MARL to tailor essential patterns and then generate a personalized prompt at sentence level, enhancing search efficiency and ensuring prompt quality.
  • Experimental results validate that RPP/RPP+ enables LLMs to outperform diverse baseline methods in ranking tasks, demonstrating the potential of instance-wise prompting.
@@ 2.1. Task-wise Prompting
Most LLM-based recommender systems frame recommendation as a language modeling task (Wang and Lim, 2023; Gao et al., 2023; Dai et al., 2023; Kang et al., 2023), with a fixed prompt template shared across all users in a task. The fixed task-wise prompts typically consist of four patterns, including the “role-playing”, “history records”, “reasoning guidance”, and “output format” patterns, as exemplified as follows:
@@ 3. Methodology
Figure 2. The framework of our proposed RPP, which can personalize instance-wise prompts for individual users via MARL. The MARL makes policies to optimize the essential patterns in a prompt respectively, maximizing the reward after interacting with the LLM environment. The “Refine” block is realized by LLMs for RPP+ to enhance the flexibility and quality of the selected actions.
To enhance LLMs’ recommendation capability with effective prompts, we introduce RPP/RPP+ to personalize instance-wise prompts with MARL, where RPP+ enhances RPP by incorporating a “refine” block. In this section, we begin by formulating prompt personalization as a Markov Decision Process, followed by a detailed description of RPP/RPP+’s components, including the action space, state space, reward function, and policy architecture.
@@ 3.1. Formulation of RL for Prompt Personalization
  • $R$ is the reward function $r(\mathbf{s}_{t},a_{t})$, which determines the reward $r\in R$ for agent after interacting with the LLMs. To integrate with recommendation tasks, we define the reward $r$ as the recommendation performance evaluated according to LLMs’ answers.
@@
For a user $u$, we initialize prompt $p_{0}$ with sentences randomly selected from each pattern (i.e., role-playing, history records, reasoning guidance, and output format). As depicted in Figure 2, we employ a multi-agent RL (MARL) under the Centralized Training with Decentralized Execution (CTDE) paradigm (Lowe et al., 2017). Each agent $z_{k}$ is responsible for personalizing a specific pattern with individual action space $\mathcal{A}_{k}$, where $k\in\{1,2,...,K\}$, $K$ is the number of agents (or patterns). Prompt personalization can be formulated as seeking policies $\pi_{k}(a_{t}^{(k)}|\mathbf{s}_{t})$ that selects actions $a_{t}^{(k)}$ from the action set $\mathcal{A}_{k}$ based on the global state $\mathbf{s}_{t}$, generating a prompt $p_{t}=concat(a_{t}^{(1)},\ldots,a_{t}^{(K)})$ at step $t$. Formally, the objective can be defined as:
@@ 3.2. Action Space
History records: The user’s historical interactions constitute critical contextual information for eliciting high-quality recommendation (Liu et al., 2023; Hou et al., 2024). Intuitively, certain users exhibit short-term interests best captured by recent interactions, while others having long-term preferences require a longer sequence. Therefore, we personalize history records by varying the interaction sequence length to match users’ dynamic preferences. Specifically, with an initial length $l_{0}$, at each step $t$, the agent’s action is deciding the number of additional interactions $\hat{l}_{t}$ to introduce in chronological order, that is, providing $l_{t}=l_{t-1}+\hat{l}_{t}$ historical context from near to far in time. By formulating the sequence expansion as action, the agent can learn an adaptive policy to tail each user’s sequence to the short versus long-term interests.
@@
We should emphasize that the presented action set serves as a representative example of personalized prompt optimization in movie recommendations. However, this framework can be readily generalized to various recommendation tasks (e.g., rating prediction, item ranking) and domains (e.g., games, music) through the substitution of task requirements or domain-specific terminology. Moreover, the current action set can be expanded with additional choices belonging to each pattern. These demonstrate the flexibility and scalability of the proposed framework.
@@ 3.3. State Space
Initializing state $\mathbf{s}_{0}$: To account for user personality and distinguish between users, for each user, we initialize the state $\mathbf{s}_{0}$ by incorporating his/her personalized features. Here we set the user-specific features as user embedding derived from the traditional recommender models (e.g., LightGCN (He et al., 2020)), which have demonstrated a powerful ability to capture collaborative signals. Formally, given the ID information of user $u$ and item $i$, we obtain their embeddings:
@@
where $\mathbf{u}$ and $\mathbf{i}$ denote the representation of user $u$ and item $i$, respectively, $f(\cdot)$ represents the encoder of traditional recommender models.

### v2

@@ Abstract.
Designing effective prompts can empower LLMs to understand user preferences and provide recommendations with intent comprehension and knowledge utilization capabilities. Nevertheless, recent studies predominantly concentrate on task-wise prompting, developing fixed prompt templates shared across all users in a given recommendation task (e.g., rating or ranking). Although convenient, task-wise prompting overlooks individual user differences, leading to inaccurate analysis of user interests. In this work, we introduce the concept of instance-wise prompting, aiming at personalizing discrete prompts for individual users. Toward this end, we propose Reinforced Prompt Personalization (RPP) to realize it automatically. To improve efficiency and quality, RPP personalizes prompts at the sentence level rather than searching in the vast vocabulary word-by-word. Specifically, RPP breaks down the prompt into four patterns, tailoring patterns based on multi-agent and combining them. Then the personalized prompts interact with LLMs (environment) iteratively, to boost LLMs’ recommending performance (reward). In addition to RPP, to improve the scalability of action space, our proposal of RPP+ dynamically refines the selected actions with LLMs throughout the iterative process. Extensive experiments on various datasets demonstrate the superiority of RPP/RPP+ over traditional recommender models, few-shot methods, and other prompt-based methods, underscoring the significance of instance-wise prompting in LLMs for recommendation. Our code is available at https://github.com/maowenyu-11/RPP.
@@
* corresponding author
copyright: acmlicensedjournal: TOISjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3716320conference: ACM Transactions on Information Systems; July 03–05, 2024; Woodstock, NYccs: Computing methodologies Discrete space searchccs: Computing methodologies Multi-agent reinforcement learningccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RSs) are widely applied in real life to offer personalized item recommendations that align with user preference. Recently, large language models (LLMs), such as ChatGPT, GPT4, PaLM-E, and LLaMA, have demonstrated exceptional capabilities in semantic understanding, intent reasoning, and knowledge utilization (Zhao et al., 2023). These advancements empower LLMs to address recommendation tasks by modeling users’ behavior into natural language, attracting considerable attention (Li et al., 2023; Geng et al., 2022; Ding et al., 2024).
To unleash the recommendation capabilities of LLMs, a promising solution lies in tailoring prompts for the specific recommendation task (Zhang et al., 2023b; Geng et al., 2022), beyond fine-tuning the massive parameters (Bao et al., 2023) that can be resource-hungry. Existing methods (Wu et al., 2023b; Gao et al., 2023; Hou et al., 2024) have investigated the task-wise prompting, dedicated to establishing fixed prompt templates shared across all users. To summarize, most of these task-wise prompts for recommendation tasks include the following patterns:
@@
  • History records provides the history interaction sequence of the target user, which the LLMs can utilize as in-context information for subsequent recommendations (Hou et al., 2024; Liao et al., 2024).
  • Reasoning guidance guides LLMs with essential steps in the reasoning process to make recommendations, incorporating strategies such as chain-of-thought (CoT) (Wei et al., 2022; Hou et al., 2024).
@@
However, we argue that applying fixed task-wise prompts for all users (Jin et al., 2023) may fail to fully explore the capability of LLMs to provide personalized recommendations. First, users’ varied intentions pose challenges for the one-fitting-all prompting to analyze dynamically, which may result in mediocre recommendation performance. Take Figure 1 as an example, the first user’s preference for the science fiction film “The Fly” is reflected in the watching history of 10 movies from two weeks ago. In contrast, the second user’s preference for the comedy film “Pink Flamingos” is evident in the watching history of the recent two films. Thus, task-wise prompts, such as instructing LLMs to recommend based on a fixed length of users’ watching history may neglect users’ dynamic preferences. Secondly, the performance of LLMs is sensitive to the expression of prompts (Lu et al., 2022). Fixed prompt templates sacrifice the benefits that varying expressions of prompts can bring to the output of LLMs. Therefore, we propose personalizing instance-wise prompts for different users in the recommendation tasks.
While conceptually appealing, personalizing the discrete prompts for individual users is challenging. On the one hand, the methods of manual creation (Brown et al., 2020) or heuristical crafting (Jiang et al., 2020)) are labor-intensive and resource-expensive to tailor prompts for each user. On the other hand, methods based on supervised learning (Rubin et al., 2022; Liu et al., 2022a) are unreliable due to the absence of labels for optimal prompts. To bridge this gap, we propose personalizing prompts automatically with reinforcement learning (RL), drawing inspiration from recent advances in optimizing prompts for natural language processing (NLP) tasks (Deng et al., 2022; Zhang et al., 2023a). The advantage of RL (Wang et al., 2024b) lies in its ability to learn autonomously through interactions with the environment, rather than depending on manual effort or labeled data. However, existing works which optimize prompts with RL in NLP tasks either search for tokens within an extensive vocabulary (Deng et al., 2022) or conduct basic editing operations (Zhang et al., 2023a) on the prompts, which are constrained by computational resource limitations and the reliability of the generated prompts. To apply RL for personalizing instance-wise prompts in the recommendation tasks, two fundamental problems are crucial.
Figure 1. Comparison between task-wise prompting and instance-wise prompting for recommendation. The ground truth for candidate movies that the users will watch next is marked in red. The personalized parts in prompts for different users are highlighted with yellow.
  • How to reduce the search space to enhance efficiency?
@@
To tackle these problems, we introduce Reinforced Prompt Personalization (RPP) to personalize high-quality prompts for each user efficiently. Instead of searching word-by-word in the extensive vocabulary (Deng et al., 2022), we limit the search space by exploring the various expressions of the four patterns at the sentence level, i.e., {role-playing, history records, reasoning guidance, output format}. We view the prompt optimization process as a combinatorial optimization problem with limited search space (Liu et al., 2025, 2024b; Kuang et al., 2024; Liu et al., 2024a). To ensure prompt quality, we meticulously design the diverse expressions of each pattern based on the nature of recommendation tasks, which can guide LLMs to understand users’ intent and make recommendations from multiple perspectives. We specify these high-quality expressions as actions, in contrast to simply treating different editing operations (such as “add” or “delete”) as the actions (Zhang et al., 2023a). The personalized prompts interact with LLMs which function as the environment, to maximize LLMs’ recommending performance which serves as the reward. Specifically, to select sentences (actions) for the four patterns with specific expertise respectively, RPP is designed based on multi-agent RL (MARL) under Centralized Training with Decentralized Execution (CTDE) paradigm  (Lowe et al., 2017). Under CTDE, each agent has an individual policy network based on Advantage Actor-Critic  (Peng et al., 2018) algorithm, as personalizing each specific prompt pattern requires distinct expertise. This global state contains information about the personalized prompts and the ranking results generated by LLMs in the current round. The action space is carefully designed with each action collected through collaborative deliberation between human experts and LLMs. Moreover, to improve the flexibility and quality of the action space, we dynamically refine the selected actions with LLMs during the iterative process to enhance scalability, referred to as RPP+.
@@
We assess the effectiveness of our approach in ranking tasks across three widely used public datasets(i.e., MovieLens-1M (Harper and Konstan, 2016), Games (Ni et al., 2019), and Lastfm (Cantador et al., 2011)). Experimental results demonstrate that RPP/RPP+ can significantly improve the recommendation performance of LLMs, outperforming several traditional recommender models (e.g., BPRMF (Rendle et al., 2009), LightGCN (He et al., 2020), SASRec (Kang and McAuley, 2018)), few-shot methods (e.g., VQ-Rec (Hou et al., 2023), UniSRec (Hou et al., 2022)), and other prompt-based methods (e.g., manual prompts (Hou et al., 2024), enumeration, and GRIPS (Prasad et al., 2023)). This highlights the significance of tailoring prompts for each individual user and the superiority of RPP/RPP+ in enhancing the recommendation ability of LLMs. Our key contributions are as follows:
@@
  • We propose Reinforced Prompt Personalization (RPP/RPP+), a general prompt optimization framework to tailor instance-wise prompts for individual users.
  • Our methodology generates personalized prompts at the sentence level with MARL, achieving the balance between search efficiency and prompt quality.
  • Extensive experiments on three benchmark datasets demonstrate the superiority of RPP/RPP+, highlighting the necessity of personalizing instance-wise prompts.
@@ 2.1. Task-wise Prompting
Most LLM-based recommender systems frame recommendation as a language modeling task (Wang and Lim, 2023; Gao et al., 2023; Dai et al., 2023), with a fixed prompt template shared across all users in a task. The fixed task-wise prompts typically consist of four patterns, including “role-playing”, “history records”, “reasoning guidance”, and “output format”, as exemplified as follows:
@@ 3. Methodology
Figure 2. The framework of our proposed RPP/RPP+. MARL serves as the core component to personalize instance-wise prompts with four distinct patterns, using corresponding agents. It is trained iteratively to maximize rewards based on the outputs of the frozen LLM-based recommender. Once trained, MARL can select optimal actions from four patterns to generate personalized prompts for each user based on their data, effectively prompting the LLM-based recommender for tailored recommendations. In addition to RPP, the “Refine” block is designed for RPP+ to enhance the flexibility and quality of the selected actions, utilizing other LLMs to dynamically refine the selected actions before prompting the LLM-based recommender.
To enhance LLMs’ recommendation capability with effective prompts, we introduce RPP/RPP+ to personalize instance-wise prompts with MARL, where RPP+ enhances RPP by incorporating a “refine” block. Comprising four actor-critic networks, the MARL serves as the core component to select actions from the four patterns and create personalized instance-wise prompts for different users. These optimized prompts are then used to prompt an LLM-based recommender for recommendations, while the MARL is trained iteratively to maximize rewards from the outputs of LLM, as shown in Figure 2. In this section, we begin by formulating prompt personalization as a Markov Decision Process, followed by a detailed description of RPP/RPP+’s components, including the action space, state space, reward function, and policy architecture.
@@ 3.1. Formulation of RL for Prompt Personalization
  • $R$ is the reward function, $r(\mathbf{s}_{t},a_{t})$ determines the reward $r\in R$ for agent after interacting with the LLMs. To integrate with recommendation tasks, we define the reward $r$ as the recommendation performance evaluated according to LLMs’ answers.
@@
For a user $u$, we initialize prompt $p_{0}$ with sentences randomly selected from $K$ patterns (i.e., role-playing, history records, reasoning guidance, and output format). As depicted in Figure 2, we employ a multi-agent RL (MARL) under the Centralized Training with Decentralized Execution (CTDE) paradigm (Lowe et al., 2017). Each agent $z_{k}$ is responsible for personalizing a specific pattern with individual action space $\mathcal{A}_{k}$, where $k\in\{1,2,...,K\}$, $K$ is the number of agents (or patterns). Prompt personalization can be formulated as seeking policies $\pi_{k}(a_{t}^{(k)}|\mathbf{s}_{t})$ that selects actions $a_{t}^{(k)}$ from the action set $\mathcal{A}_{k}$ based on the global state $\mathbf{s}_{t}$, generating a prompt $p_{t}=concat(a_{t}^{(1)},\ldots,a_{t}^{(K)})$ at step $t$. The generated prompt $p_{t}$ interacts with LLMs (environments) to enhance LLMs’ recommendation performance (reward) by maximizing the objective of our multi-agent. Formally, the objective can be defined as:
@@ 3.2. Action Space
History records: The user’s historical interactions constitute critical contextual information for eliciting high-quality recommendation (Liu et al., 2023b; Hou et al., 2024). Intuitively, certain users exhibit short-term interests best captured by recent interactions, while others having long-term preferences require a longer interaction history. Therefore, we personalize history records by varying the interaction sequence length to match users’ dynamic preferences. Specifically, with an initial length $l_{0}$, at each step $t$, the agent’s action is deciding the number of additional interactions $\hat{l}_{t}$ to introduce in chronological order, that is, providing $l_{t}=l_{t-1}+\hat{l}_{t}$ historical context from near to far in time. By formulating the sequence expansion as action, the agent can learn an adaptive policy to tail each user’s sequence to the short versus long-term interests.
@@
We should emphasize that the presented action set serves as a representative example of personalized prompt optimization in movie recommendations. However, this framework can be generalized to various recommendation tasks (e.g., rating prediction, item ranking) and domains (e.g., games, music) through the substitution of task requirements or domain-specific terminology. Moreover, the current action set can be expanded with additional choices belonging to each pattern. These demonstrate the flexibility and scalability of the proposed framework.
Refine actions in RPP+: In addition to RPP, the “refine” block designed for RPP+ can dynamically refine the selected actions with other LLMs before prompting the LLM-based recommenders during each iteration, which can automatically enhance the flexibility and quality of the selected actions. To refine the selected sentence action with LLMs, we use the prompt as “Please refine this sentence to effectively prompt LLMs for recommendations”. After the “refine” block of RPP+, we obtain a more polished sentence for prompting compared to that generated by RPP.
@@ 3.3. State Space
Initializing state $\mathbf{s}_{0}$: Since state information after decisions is unavailable initially, we initialize the state $\mathbf{s}_{0}$ by incorporating users’ personalized features to differentiate between them. Here we set the user-specific features as user embedding derived from the traditional recommender models (e.g., LightGCN (He et al., 2020)), which can capture collaborative signals with graphs (Wu et al., 2023a, 2022). Formally, given the ID information of user $u$ and item $i$, we obtain their embeddings:
@@
where $\mathbf{u}$ and $\mathbf{i}$ denote the representation of user $u$ and item $i$, respectively, $f(\cdot)$ represents the encoder of traditional recommender models. Based on the personalized information of each user, RPP/RPP+ makes policy to obtain the initial actions $a_{0}^{(k)}$ and initial prompts $p_{0}$.